# MVP-ORAM: a Wait-free Concurrent ORAM for Confidential BFT Storage

### v1

@@ VI Byzantine Fault-Tolerant ORAM
The previous section detailed MVP-ORAM, a protocol that can handle concurrent clients accessing an ORAM while satisfying wait freedom and Linearizability. We now describe how MVP-ORAM can be replicated using BFT SMR to tolerate fully malicious servers, ensuring data integrity and availability while preserving data and access-pattern secrecy. For this, we use a Byzantine Fault-Tolerant State Machine Replication (BFT SMR) protocol [castro1999practical, schneider1990implementing, bessani_2014].
BFT SMR is a classical technique for implementing fault-tolerant systems by replicating stateful, deterministic services on multiple fault-independent servers [schneider1990implementing]. Most BFT SMR implementations allow tolerating $t$ Byzantine faults with $n>3t$ servers. This is possible by making each server starts in the same initial state and executes the same sequence of operations deterministically. Ensuring such a total order of operations on all correct servers requires executing Byzantine consensus [castro1999practical, yin_2019] to make the replicas agree on the next set of client operations to be executed.
MVP-ORAM solves three fundamental challenges that are required for replicating an ORAM through BFT SMR. First, it ensures the server-side algorithm is fully deterministic. Second, it requires only three invocations of state machine operations ($\mathsf{getPM}$, $\mathsf{getPS}$, and $\mathsf{evict}$) for performing an access. Third, and most importantly, it makes these operations wait-free.
In detail, we execute $n$ server replicas using a BFT SMR middleware to ensure that the invocation of the three server operations used in Algorithm 1 is reliably disseminated in total order to all servers. Each server executes those functions locally, exactly as specified in Algorithm 2, and sends replies to the invoking clients, which consolidate a single response for each invocation by waiting for $t+1$ matching replies.
@@
The first optimization is related to the execution of consensus over metadata. Byzantine consensus protocols typically select one (the leader, as in PBFT [castro1999practical]) or more (the DAG block proposers, as in Mysticeti [babel2025mysticeti]) proposers to disseminate batches of requests to be ordered. In such protocols, the client sends its request to all replicas, and proposers re-disseminates the request along with ordering information. However, if the clients’ requests are large (as in our case, where $\mathsf{evict}$ has to send a path, path map, and stash), the proposer bandwidth will be easily exhausted. To solve this, the client can send the $\mathsf{evict}$ parameters directly to the servers without ordering them and send only their hash for ordering through BFT SMR. Using the hashes, replicas retrieve the evict’s data and process the request as usual, thus significantly reducing the best-case bandwidth usage.
@@ VII-A Security Analysis
When a sequence of requests $\overrightarrow{y}$ is performed, the servers see $A(\overrightarrow{y})$, which is the same sequence of requests but transformed by the ORAM. When $c$ clients access different blocks within the same timestep (case 2), each of them selects a random leaf, resembling the behavior of a single client performing $c$ accesses across $c$ consecutive timesteps (case 1). In Lemma 7, we show that in these first two cases, $A(\overrightarrow{y})$ becomes indistinguishable from a random sequence of requests with high probability.
@@
Fig. 5 shows the statistical distance of the distributions (a worst-case measurement), considering different values of $L$ and $\alpha$. As expected, the distance decreases as we make the accesses more skewed (i.e., as we increase $\alpha$) and as we increase the tree size $L$. If we decrease $\alpha$ to a point where the number of frequently accessed blocks approximates $N$ (i.e., blocks are uniformly accessed), the statistical distance will be near $1$ (worst security). It is worth to remark that although statistical security is not good, the concrete probabilities of leakage are very small. The expected probability of $c$ clients accessing the same block using the same path with uniform access distribution is $\frac{1}{2N^{c-1}}$. For example, this probability for two clients and a tree of $N=2^{18}$ is less than $0.0002\%$.
@@ VII-C Bandwidth and Storage Analysis
We now analyze the communication complexity of MVP-ORAM. To simplify this analysis, we omit the cost of sending constant values such as block and client ids. We start by considering a single server (non-BFT) setup. The $\mathsf{getPM}$ reply contains $c$ path maps of size $O(c+\log N)$, resulting in $O(c(c+\log N))$ bits. The $\mathsf{getPS}$ reply has size $O(c^{2}\log N)$, as it contains $c$ paths and stashes. Finally, an $\mathsf{evict}$ has size $O(c\log N)$ (a consolidated path and stash).
When considering a $n$-server setup, by default, the communication goes up by at least a factor of $n$, as data has to be sent to all servers. Notice this multiplicative factor only happens in requests, as replies benefit from the optimization of n1n-1 servers sending hashes, i.e., $O_{\mathit{reply}}=O(|\mathit{Reply}|+n)$. Requests also account for the cost of Byzantine consensus, which we use only for metadata ordering, which results in $O_{\mathit{request}}=O(n|\mathit{Request}|+\mathit{Consensus})$.
@@ VIII-A Mitigating the Risk of Concurrent Accesses
For this strategy to work, concurrent clients need to define different timesteps for their real accesses. A client $c_{i}$ discovers $\tau_{i}$ during its first MVP-ORAM access. In the invocation of $\mathit{getPM}$, each client additionally sends the (encrypted) $\mathit{addr}$ that it wants to access, and the server stores this information in a set $\mathcal{A}$, which is returned together with the $\mathit{getPM}$’s response.
@@ VIII-B Security Analysis of Strong MVP-ORAM
Our strategy requires every client to always perform $\sigma+1$ accesses, ensuring in case concurrent clients access the same address, they do so at different timesteps. Since the server does not know which address a client is accessing, the adversary will be unable to distinguish between real and dummy accesses and whether clients are coordinating to hide access to the same address. Hence, as long as $\sigma$ matches the maximum concurrency of the system ($c\leq\sigma+1$), then this approach ensures no two clients access the same block in the same timestep and it is possible to show that the access pattern will be indistinguishable from a random access pattern, matching the obliviousness of parallel ORAM without giving up wait-freedom but assuming synchrony.
@@ IX-A Implementation
We built a prototype of MVP-ORAM in Java by extending COBRA [vassantlal_2022], a confidential BFT SMR framework based on DPSS. COBRA itself relies on BFT-SMaRt [bessani_2014], a replication library providing all features required for practical BFT SMR systems. BFT-SMaRt implements a Verifiable and Provable Consensus [sousa_2012] based on Cachin’s Byzantine Paxos [cachin_2009], which is similar to PBFT [castro1999practical], i.e., it requires three communication steps and has a quadratic message complexity. This is the consensus algorithm executed during the invocation of $\mathsf{Server.getPM}$, $\mathsf{Server.getPS}$, and $\mathsf{Server.evict}$. Thus, through COBRA (and BFT-SMaRt), we can easily implement all features required by MVP-ORAM.
@@
Our implementation and all the code used for the experiments in this paper is available on the project’s web page [mvp_oram].
@@ IX-B Setup and Methodology
Our experimental evaluation was performed in the AWS cloud using two types of instances. The servers were executed in $n$ r5n.2xlarge instances, each having $8$ vCPU, $64$ GB of RAM, and $8.1$ Gbps baseline bandwidth. The clients were executed in $6$ c5n.2xlarge instances, each with $8$ vCPU, $21$ GB of RAM, and $10$ Gbps baseline bandwidth.
We installed Ubuntu Server 22.04 LTS and OpenJDK 11 on all of the machines. The experiment analyzing stash size was conducted by simulating concurrent accesses to ORAM on a single machine. Performance was measured by executing servers in $n$ machines and clients in the remaining ones. Throughput and latency measurements were collected from a single server and client machine, respectively. Unless stated otherwise, we measure MVP-ORAM considering a database of $1$ GB configured in a tree of height $L=17$ and with each node containing $Z=4$ blocks of $4096$ bytes each. Therefore, our database contains $N=262143$ blocks. For each operation, clients randomly choose among all the blocks in a Zipfian distribution with $\alpha=1.0$ (unless stated otherwise), and the operation type is picked uniformly at random between read and write. The graphs showing throughput and latency display the average (plus standard deviation) of data collected over $6$ minutes for each experiment.
@@ IX-C MVP-ORAM Stash Size
Since altering $\alpha$ affects the number of frequently accessed addresses and $c$ affects concurrency, this experiment analyses the impact of changing those parameters. The bucket size $Z$ also affects the stash usage. However, it has the expected result: the stash size decreases as we increase $Z$ since the probability of selecting empty slots from the path increases, which increases the number of blocks evicted from the stash.
@@
Figure 7: Stash size for different Zipfian’s exponent values ($\alpha$), with $c=15$, and number of clients ($c$), with $\alpha=1.0$. Each point is an average of 5050k accesses.
As analyzed in §VII-B, the rate of moving blocks from the stash to the tree and vice-versa is equal when the stash size reaches its maximum size.
We study the impact of Zipfian exponent $\alpha$ by experimenting with low ($\alpha=10^{-6}$), medium ($\alpha=1.0$), and high ($\alpha=2.0$) contention levels in the accessed blocks. Accordingly, clients approximately access $90\%$, $27\%$, and $0.001\%$ of blocks, respectively, with a probability more than $90\%$. Decreasing this parameter increases the number of distinct blocks accessed and moved to the stash by clients. However, the number of concurrent clients heavily influences the number of distinct blocks swapped between the stash and the tree. Since $c$ is fixed, the stash size is defined by the number of frequently accessed blocks, which is higher for small $\alpha$ values.
Regarding the number of concurrent clients, increasing that parameter also increases the stash size. Since blocks are uniformly sampled, increasing the number of clients increases the number of common blocks selected from the stash. Thus, the clients remove a few distinct blocks from the stash, increasing the stash size. This experiment shows that the number of concurrent clients heavily dominates the stash size, confirming our theoretical analysis (§VII-B).
@@ IX-D MVP-ORAM Performance
The overall throughput of the single-server setup is higher than that of the system tolerating failures. Specifically, the peak throughput is $2.2\times$ the system’s throughput tolerating one failure. Recall that a system tolerating $t$ faults requires $n>3t$ servers, and clients must wait for at least $t+1$ matching server responses before continuing. Therefore, as the number of tolerated faults increases, clients must send requests to multiple servers and wait for their responses, which increases latency and reduces the number of requests they send. In turn, this slightly reduces the overall throughput, i.e., it drops by $7\%$ when the number of servers goes from $4$ to $10$.
@@
To better understand the factors contributing to MVP-ORAM’s performance, we break down the access latency in the three operations. Fig. 9 shows the individual latency of each protocol phase, which corresponds to an SMR operation for different numbers of servers and up to $c_{max}$ clients, when queueing is not a factor. With a single client (no concurrency - first bar of each group), $\mathsf{evict}$ is the most costly phase. When the number of concurrent clients increases (second and third bars), $\mathsf{getPS}$ becomes more prominent. This can be explained by the operations’ request/response sizes for different clients, as shown in Table III. Among the three phases of MVP-ORAM, $\mathsf{getPS}$ requires bandwidth that is quadratic in the number of clients (see Table II).
@@
TABLE III: MVP-ORAM access request/response sizes in bytes, for workloads with 11, 55, and $10$ clients.
@@ 55 $8/1489$ $8/445932$ $314101/1$
Strong MVP-ORAM performance. The strong variant of MVP-ORAM discussed in §VIII requires each ORAM access to perform $\sigma+1$ MVP-ORAM accesses. This means the throughput observed for this variant would be MVP-ORAM throughput divided by $\sigma+1$. For example, if one wants to offer perfect ORAM guarantees for up to $10$ clients, each access would require eleven MVP-ORAM accesses, which means a throughput of about the $\approx 390/11=35$ accesses/sec. A similar degradation also affects latency.
@@ IX-E Performance with Different Configurations
Our last set of experiments considers the performance of MVP-ORAM with different block sizes: 256 bytes, 1024 bytes, and 4096 bytes. Fig. 12 shows the throughput and latency for the experiments.
For small blocks of 256 bytes, the system reaches almost a thousand accesses per second. As the block size increases, performance decreases accordingly, since 4 and 16 times more data is transferred with the other block sizes.
@@
Overall, these additional experiments confirm our theoretical observation that MVP-ORAM performance gets worse in configurations that require more data to be transferred. This aligns with observations made for other ORAM protocols and supports the fundamental goal of enhancing the overall bandwidth of ORAM schemes.
@@ IX-F Experimental comparison with other systems
QuORAM [maiyya2022quoram] is the only replicated ORAM service we are aware of. It tolerates crash faults and uses trusted proxies, while MVP-ORAM tolerates Byzantine faults without requiring trusted components. The existence of trusted proxies collocated with servers in the same machine eliminates the need to execute bandwidth-hungry ORAM operations through the network, as they are executed only between the proxy (which acts as a single Path ORAM client) and the server. We confirmed the benefit QuORAM gained with this approach by executing it in the same setting as MVP-ORAM. With $100+$ clients and for $n=4$ and $n=7$, QuORAM achieved a throughput of around $1000$ ops/s. However, our evaluation also shows that QuORAM performs much worse than MVP-ORAM with a restricted number of clients. As sumarized in Table IV, MVP-ORAM with $50$ client can process $356$ ($n=4$) and $355$ ($n=7$) ops/s, while QuORAM only processes between $163$ and $183$ ops/s. The reason behind the low performance of QuORAM is its high initial latency, due to the use of proxies. In our experiments, MVP-ORAM’s maximum latency is $130$ ms, while QuORAM’s minimum latency is $272$ ms.
@@
We argue that having too many clients increases the service attack surface, as ORAM clients must be mutually trusted. Therefore, we argue that achieving high performance with fewer clients is more important than reaching good numbers with $100+$ concurrent clients accessing the ORAM.

### v2

@@ VI Byzantine Fault-Tolerant ORAM
The previous section detailed MVP-ORAM, a protocol that can handle concurrent clients accessing an ORAM while satisfying wait-freedom and linearizability. We now describe how MVP-ORAM can be replicated using BFT SMR to tolerate fully malicious servers, ensuring data integrity and availability while preserving data and access-pattern secrecy. For this, we use a Byzantine Fault-Tolerant State Machine Replication (BFT SMR) protocol [castro1999practical, schneider1990implementing].
BFT SMR is a classical technique for implementing fault-tolerant systems by replicating stateful, deterministic services on multiple fault-independent servers [schneider1990implementing]. Most BFT SMR implementations allow tolerating $t$ Byzantine faults with $n>3t$ servers. This is possible by ensuring that each server starts in the same initial state and executes the same sequence of operations deterministically. Ensuring such a total order of operations on all correct servers requires executing Byzantine consensus [castro1999practical, yin_2019] to make the replicas agree on the next set of client operations to be executed.
MVP-ORAM solves three fundamental challenges that are required for replicating an ORAM through BFT SMR. First, it ensures the server-side algorithm is fully deterministic. Second, it requires only three invocations of state machine operations ($\mathsf{getPM}$, $\mathsf{getPS}$, and $\mathsf{evict}$) for performing an access. Third, and most importantly, it makes ORAM accesses wait-free.
In detail, we execute $n$ server replicas using a BFT SMR middleware (e.g., [bessani_2014]) to ensure that the invocation of the three server operations used in Algorithm 1 is reliably disseminated in total order to all servers. Each server executes those functions locally, exactly as specified in Algorithm 2, and sends replies to the invoking clients, which consolidate a single response for each invocation by waiting for $t+1$ matching replies.
@@
The first optimization is related to the execution of consensus over metadata. Byzantine consensus protocols typically select one (the leader, as in PBFT [castro1999practical]) or more (the DAG block proposers, as in Mysticeti [babel2025mysticeti]) proposers to disseminate batches of requests to be ordered. In such protocols, the client sends its request to all replicas, and proposers re-disseminate the request along with ordering information. However, if the clients’ requests are large (as in our case, where $\mathsf{evict}$ must send a path, path map, and stash), the proposer’s bandwidth will be easily exhausted. To solve this, the client can send the $\mathsf{evict}$ parameters directly to the servers without ordering them and send only their hash for ordering through BFT SMR. Using the hashes, replicas retrieve the operation parameters and process the request as usual, thus significantly reducing the best-case bandwidth usage.
@@ VII Security and Complexity Analysis
We now discuss the security, correctness, and theoretical performance of MVP-ORAM.
@@ VII-A Security Analysis
When a sequence of requests $\overrightarrow{y}$ is performed, the servers see $A(\overrightarrow{y})$, which is the same sequence of requests but transformed by the ORAM. When $c$ clients access different blocks within the same timestep (case 2), each of them selects a random leaf, resembling the behavior of a single client performing $c$ accesses across $c$ consecutive timesteps (case 1). In Lemma 7 we show that $A(\overrightarrow{y})$ becomes indistinguishable from a random sequence of requests with high probability in case 1 (case 2 is ommitted, since they are similar).
@@
Fig. 5 shows the statistical distance of the distributions (a worst-case measurement), considering different values of $L$ and $\alpha$. As expected, the distance decreases as we make the accesses more skewed (i.e., as we increase $\alpha$) and as we increase the tree size $L$. If we decrease $\alpha$ to a point where the number of frequently accessed blocks approximates $N$ (i.e., blocks are uniformly accessed), the statistical distance will be near $1$ (worst security). Nonetheless, it is worth noting that although the statistical security in this case is far from good, the concrete probabilities of leakage are very small. The expected probability of $c$ clients accessing the same block using the same path with uniform access distribution is $\frac{1}{2N^{c-1}}$. For example, this probability for two clients and a tree of $N=2^{18}$ is less than $0.0002\%$.
@@ VII-C Bandwidth and Storage Analysis
We now analyze the communication complexity of MVP-ORAM. To simplify this analysis, we omit the cost of sending constant values such as block and client ids. We begin by considering a single-server (non-BFT) setup. The $\mathsf{getPM}$ reply contains up to $c$ path maps of size $O(c+\log N)$, resulting in $O(c(c+\log N))$ bits. The $\mathsf{getPS}$ reply has size $O(c^{2}\log N)$, as it contains $c$ paths and stashes. Finally, an $\mathsf{evict}$ has size $O(c\log N)$ (a consolidated path and stash).
When considering an $n$-server setup, by default, the communication goes up by at least a factor of $n$, as data must be sent to all servers. Note that this multiplicative factor only occurs in requests, as replies benefit from the optimization of n1n-1 servers sending hashes; i.e., $O(|\mathit{Reply}|+n)$. Requests also account for the cost of Byzantine consensus, which we use only for metadata ordering, which results in $O(n|\mathit{Request}|+\mathit{Consensus})$.
@@ VIII-A Mitigating the Risk of Concurrent Accesses
For this strategy to be effective, concurrent clients must define distinct timesteps for their actual accesses. A client $c_{i}$ discovers $\tau_{i}$ during its first MVP-ORAM access. In the invocation of $\mathsf{getPM}$, each client additionally sends the (encrypted) $\mathit{addr}$ that it wants to access, and the server stores this information in a set $\mathcal{A}$, which is returned together with the $\mathsf{getPM}$’s response.
@@ VIII-B Security Analysis of Strong MVP-ORAM
Our strategy requires every client to always perform $\sigma+1$ accesses, ensuring that in case concurrent clients access the same address, they do so at different timesteps. Since the server does not know which address a client is accessing, the adversary will be unable to distinguish between real and dummy accesses and whether clients are coordinating to hide access to the same address. Hence, as long as $\sigma$ matches the maximum concurrency of the system ($c\leq\sigma+1$), then this approach ensures no two clients access the same block in the same timestep and it is possible to show that the access pattern will be indistinguishable from a random access pattern, matching the obliviousness of parallel ORAM without giving up wait-freedom but assuming synchrony.
@@ IX-A Implementation
We built a prototype of MVP-ORAM in Java by extending COBRA [vassantlal_2022], a confidential BFT SMR framework based on DPSS. COBRA itself relies on BFT-SMaRt [bessani_2014], a replication library that provides all the features required for practical BFT SMR systems. BFT-SMaRt implements a Verifiable and Provable Consensus [sousa_2012] based on Cachin’s Byzantine Paxos [cachin_2009], which is similar to PBFT [castro1999practical], i.e., it requires three communication steps and has a quadratic message complexity in the common case. This is the consensus algorithm executed during the invocation of $\mathsf{Server.getPM}$, $\mathsf{Server.getPS}$, and $\mathsf{Server.evict}$. Thus, through COBRA (and BFT-SMaRt), we can easily implement all features required by MVP-ORAM.
@@
Our implementation and all the code used for the experiments are available on the project’s web page [mvp_oram].
@@ IX-B Setup and Methodology
Our experimental evaluation was performed in the AWS cloud using two types of instances. The servers were executed in $n$ r5n.2xlarge instances, each having $8$ vCPU, $64$ GB of RAM, and $8.1$ Gbps baseline network bandwidth. The clients were executed in $6$ c5n.2xlarge instances, each with $8$ vCPU, $21$ GB of RAM, and $10$ Gbps baseline network bandwidth.
We installed Ubuntu Server 22.04 LTS and OpenJDK 11 on all of the machines. The experiment analyzing stash size was conducted by simulating concurrent accesses to ORAM on a single machine. Performance was measured by executing servers in $n$ machines and clients in the remaining ones. Throughput and latency measurements were collected from a single server and client machine, respectively. Unless stated otherwise, we measure MVP-ORAM considering a database of $1$ GB configured in a tree of height $L=17$ and with each node containing $Z=4$ blocks of $4096$ bytes each. Therefore, our database contains $N=262143$ blocks. For each operation, clients randomly choose among all the blocks in a Zipfian distribution with $\alpha=1.0$, and the operation type is picked uniformly at random between read and write. The graphs showing throughput and latency display the average (plus standard deviation) of data collected over $6$ minutes for each experiment.
@@ IX-C MVP-ORAM Stash Size
Since altering $\alpha$ affects the number of frequently accessed addresses and $c$ affects concurrency, this experiment analyzes the impact of changing those parameters. The bucket size $Z$ also affects the stash usage. However, it has the expected result: the stash size decreases as we increase $Z$ since the probability of selecting empty slots from the path increases, which increases the number of blocks evicted from the stash.
@@
Figure 7: Stash size with $c=15$ and different Zipfian’s exponent values, and $\alpha=1.0$ and different number of clients. Each point represents an average of 50k accesses.
As analyzed in §VII-B, the rate of moving blocks from the stash to the tree and vice-versa is equal when the stash size reaches its expected maximum size.
We study the impact of Zipfian exponent $\alpha$ by experimenting with low ($\alpha=10^{-6}$), medium ($\alpha=1.0$), and high ($\alpha=2.0$) contention levels in the accessed blocks. Accordingly, clients approximately access $90\%$, $27\%$, and $0.001\%$ of blocks, respectively, with a probability greater than $90\%$. Decreasing this parameter increases the number of distinct blocks that clients access and move to the stash. However, the number of concurrent clients heavily influences the number of distinct blocks swapped between the stash and the tree. Since $c$ is fixed, the stash size is defined by the number of frequently accessed blocks, which is higher for small $\alpha$ values.
Increasing the number of concurrent clients also increases the stash size. Since blocks are uniformly sampled, increasing the number of clients increases the number of common blocks selected from the stash. Thus, the clients remove a few distinct blocks from the stash, increasing its size. This experiment shows that the number of concurrent clients heavily dominates the stash size, confirming our theoretical analysis (§VII-B).
@@ IX-D MVP-ORAM Performance
The overall throughput of the single-server setup is higher than that of the system tolerating failures. Specifically, the peak throughput is $2.2\times$ the system’s throughput tolerating one failure. Recall that a system tolerating $t$ faults requires $n>3t$ servers, and clients must wait for at least $t+1$ matching server responses before continuing. Therefore, as the number of tolerated faults increases, clients must send requests to more servers and wait for their responses, which increases latency and reduces the number of requests they send. In turn, this slightly reduces the overall throughput, i.e., it drops by $7\%$ when the number of servers goes from $4$ to $10$.
@@
To better understand the factors contributing to MVP-ORAM’s performance, we break down the access latency in the three operations that constitute an access. Fig. 9 shows the individual latency of each protocol phase, which corresponds to an SMR operation for different numbers of servers and up to $c_{max}$ clients, when queueing is not a factor. With a single client (no concurrency - first bar of each group), $\mathsf{evict}$ is the most costly phase. When the number of concurrent clients increases (second and third bars), $\mathsf{getPS}$ becomes more prominent. This can be explained by the operations’ request/response sizes for different clients, as shown in Table III. Among the three phases of MVP-ORAM, $\mathsf{getPS}$ requires bandwidth that is quadratic in the number of clients (see Table II).
@@
TABLE III: MVP-ORAM access request/response sizes in bytes, for workloads with $1$, 55, and $10$ clients.
@@ 55 $8/1489$ $8/445932$ $314101/1$
A note on the Strong MVP-ORAM performance. The strong variant of MVP-ORAM discussed in §VIII requires each ORAM access to perform $\sigma+1$ MVP-ORAM accesses. This means the throughput observed for this variant would be MVP-ORAM throughput divided by $\sigma+1$. For example, if one wants to offer perfect ORAM guarantees for up to $10$ clients, each access would require eleven MVP-ORAM accesses, which means a throughput of about the $\approx 390/11=35$ accesses/sec. A similar degradation also affects latency.
@@ IX-E Performance with Different Configurations
Our final set of experiments evaluates the performance of MVP-ORAM with block sizes of 256 bytes, 1024 bytes, and 4096 bytes. Fig. 12 shows the throughput and latency for the experiments.
For small blocks of 256 bytes, the system reaches almost a thousand accesses per second. As the block size increases, performance decreases accordingly, since $4\times$ and $16\times$ more data is transferred with the other block sizes.
@@
Overall, these experiments confirm our theoretical observation that MVP-ORAM performance gets worse in configurations that require more data to be transferred. This aligns with observations made for other ORAM protocols and supports the fundamental goal of enhancing the overall bandwidth of ORAM schemes.
@@ IX-F Experimental comparison with other systems
QuORAM [maiyya2022quoram] is the only replicated ORAM service we are aware of. It tolerates crash faults and uses trusted proxies, while MVP-ORAM tolerates Byzantine faults without requiring trusted components. The existence of trusted proxies collocated with servers in the same machine eliminates the need to execute bandwidth-hungry ORAM operations through the network, as they are executed only between the proxy (which acts as a single Path ORAM client) and the server. We confirmed the benefit QuORAM gained with this approach by executing it in the same setting as MVP-ORAM. With over $100$ clients and for $n=4$ and $n=7$, QuORAM achieves a maximum throughput of around $1000$ operations per second. However, our evaluation also shows that QuORAM performs much worse than MVP-ORAM with a restricted number of clients. As sumarized in Table IV, MVP-ORAM with $50$ client can process $356$ ($n=4$) and $355$ ($n=7$) ops/s, while QuORAM only processes between $163$ and $183$ ops/s. The reason behind the low performance of QuORAM is its higher latency, which is caused by the use of proxies. In our experiments, MVP-ORAM’s maximum latency is $130$ ms, while QuORAM’s minimum latency is $272$ ms.
@@
Notice that having too many clients increases the service attack surface, as ORAM clients must be mutually trusted. Therefore, we argue that achieving high performance with fewer clients is more important than achieving good numbers with $100$ or more concurrent clients accessing the ORAM.